### v5

@@ 4 Case Study
Figure 3: Overall seagrass DBN model [26] (Fig. 3a). Nodes are ovals and arrows denote conditional dependence between a parent and child node in the same time-slice. Where an arrow is labelled with a $1$, the child node is in the next time-slice. Rounded rectangles denote subnetworks. Yellow nodes relate to loss and recovery in shoot density, purple nodes to recovery, green nodes to resistance, blue nodes to environmental factors and pink for all other nodes. The environment subnetwork is not shown as it is a disjoint subnetwork of the blue nodes, which are effectively inputs to the model.
@@
Table 4: Distribution of in- and out-degree (degree and number of nodes with that degree) two time-slice DBN.
@@
4.1 Computation Cost as an Objective Function
@@
Without loss of generality, consider a discrete BN. We formulate a backwards probability measure $P^{b}_{kl}$ for computing inference akin to the rearrangement of Equation (3), which forms the basis of inference techniques such as variable elimination [5].
$P^{b}_{kl}=\prod_{\forall X_{i}\in X_{kl}}\sum_{X_{i}}P\bigl{(}X_{i}|Par(X_{i})\bigr{)}\prod_{X_{kl^{\prime}}\in Child(X_{kl})}\sum_{\forall X\notin X_{kl^{\prime}}}P^{b}_{kl^{\prime}}$ (18)
$P^{b}_{kl}$ comprises propagation from child cluster-layer $P^{b}_{kl^{\prime}}$, marginalising (i.e. summing) out any nodes that are not in $\boldsymbol{X}_{kl}$, multiplication by the conditional probabilities of the nodes in $\boldsymbol{X}_{kl}$, and finally marginalisation in preparation for propagation to the next antecedent cluster-layer. It is a recursive relation and for $l^{\prime}=0$, $P^{b}_{kl^{\prime}}=1$ since a cluster-layer comprising leaf nodes does not have child nodes, by Definition 2.1. In the context of $c\Bigl{(}\boldsymbol{X}_{kl^{\prime}},u(\boldsymbol{X}_{kl^{\prime}}),Child(\boldsymbol{X}_{kl^{\prime}}),u\bigl{(}Child(\boldsymbol{X}_{kl^{\prime}})\bigr{)}\Bigr{)}$ (10), the limits of the products and sums incorporates the effect of cluster mappings $u$, and $P(X_{i}|Par(X_{i}))$ and $P^{b}_{kl^{\prime}}$ are defined with respect to cluster-layer and child cluster-layer nodes.
Define $\mathcal{F}()$ as a normalised count of the number of clock cycles associated with performing all the multiplications and additions, where each multiplication is assumed to have a normalised count of $1$ cycle, each addition $0.6$ and each division (i.e. ratio) $3$, based on Intel Skylake CPU latencies [6]. Hence,
$c\Bigl{(}\boldsymbol{X}_{kl^{\prime}},u(\boldsymbol{X}_{kl^{\prime}}),Child(\boldsymbol{X}_{kl^{\prime}}),u\bigl{(}Child(\boldsymbol{X}_{kl^{\prime}})\bigr{)}\Bigr{)}=\mathcal{F}\bigl{(}P^{b}_{kl}\bigr{)}$ (19)
4.1.1 Results and Discussion
@@
Figure 4: Cost $G$ of solutions found and the iteration of their discovery for both the one and two time-slice DBNs.
@@
Figure 5: Similarity of solutions to the set of optimal solutions and their iteration of discovery for both the one and two time-slice DBNs.
@@
5 Discussion
Clustering has been applied in a variety of approaches to improve the computational feasibility of BN learning, whether based on constraints, scoring (i.e. using a cost function) or both [10]. In this context, the vast majority of existing literature has focused on cost functions based on inclusion or exclusion of edges. Cost functions that vary with the mapping of nodes to different clusters are currently understudied. Such flexibility enables explicit optimisation of a given BN and its cluster mapping which could potentially lead to better BN learning and inference outcomes. However, finding globally optimal (i.e. least cost) cluster mappings in a BN incurs additional complexity in the form of cost dependency, where the local cost of a cluster is dependent on the mapping of connected nodes and the global cost is typically the sum of local costs.
@@ 5 Discussion
Previous BN literature has used decomposability of scoring functions, such as through additive costs, to allow for local search procedures and/or optimisation of the score function; some even use dynamic programming and graph layers [7, 12, 18]. Our work builds upon the application of dynamic programming and layers in BNs to facilitate global optimisation with arbitrary non-negative cost functions and cost dependency. This latter property has not been explicitly studied in the existing BN literature.
@@
The proposed algorithm, DCMAP, is proven to find all of the least cost cluster mapping solutions through conditional optimality of layers and branching (Section 3.2.3). In addition, it is theoretically shown that an equally informed algorithm cannot find these solutions in fewer iterations. This is important as optimal clustering is NP-hard in general [4] and cost dependency would exacerbate this complexity. For instance, a breadth first search on the same one time-slice case study is unable to find any optimal solution even with ten times more iterations than DCMAP. This arises even though the number of iterations needed is similar to DCMAP for the simple illustrative BN (Fig. 1).
5.1 Case Study Discussion
@@
We developed a case study of BN clustering where clusters were used as a means to compute (i.e. infer) the marginal probabilities of a complex systems seagrass DBN. By explicitly defining the cost function as the computational cost of inference given a cluster mapping, a speed-up of 22.8 and 426 times was achieved for a one and two time-slice version of the DBN, respectively. In itself, a substantial reduction of computational complexity through clustering is useful as BN inference plays a key role in prediction and model learning, and methods for that such as Markov Chain Monte Carlo (MCMC) [8] can require thousands of iterations. MCMC can also help better quantify uncertainty for parameter and structure learning of BNs [5] and DBNs [24] through the Bayesian framework. Explicit optimisation of BN clusters using DCMAP could be a key enabler for future research in BN learning and updating, especially for complex systems DBNs like our seagrass case study.
@@
Although our case study was for a discrete BN with between two and fourteen states per node with corresponding CPTs, the proposed DCMAP algorithm equally applies to discrete or continuous or hybrid BNs, DBNs, Object Oriented Bayesian Networks (OOBNs) [13] and Markov models in general, all of which can be represented as a DAG; the key adaptation required is the development of an application-specific cost function. Note that dynamic and templated models alike can be unrolled [26] into a BN and cluster mapping performed on the result. In practice, for a $k^{th}$ order dynamic model, running cluster mapping on $k$ slices is sufficient to find optimal clusters that can then be replicated over all slices.
@@
We have theoretically shown that an equally informed algorithm cannot find all optimal solutions in fewer iterations, further study is needed to better understand the computational complexity of DCMAP and how it might vary with different BN graph structures and cost functions. Early results from the case study suggest a complexity of $O(n^{2})$ iterations to find the first least cost cluster mapping and $O(an^{2})$ iterations for all optimal solutions, where $n$ is the number of nodes and $a$ some positive integer. This aligns with previous findings of $O(n^{2})$ complexity for clustering in the absence of cost dependency [29]. However, the computational complexity could potentially vary significantly depending on the cost function employed and also the structure of the network (e.g. dense versus sparse). This is an avenue for future research.
@@
5.2 Future Work: Entropy
In light of the generalisability of DCMAP given its ability to optimise for arbitrary, non-negative and dependent cost functions, one key area of future work is in the development of objective functions for specific usage scenarios. One such possibility is entropy, which for BNs is defined generally as the sum of the entropy associated with each scenario over all possible scenarios, where a scenario might be defined in terms of one or more nodes [23, 20]. Consider a discrete BN. The entropy associated with evidence for a given scenario indexed by $k$ is:
$H_{k}=-\sum_{k=1}^{n^{k}}p_{k}\sum_{i=1}^{n}\sum_{j=1}^{n^{s}_{i}}p_{ijk}log(p_{ijk})$ (20)
@@ 5.2 Future Work: Entropy
Consider a special case of clustering where specific evidence scenarios are to be applied at one or more nodes and at one or more points in time in a DBN, with the goal of minimising global entropy. Thus, each node can be assigned to one of two clusters: (cluster 1) observed using virtual evidence $\delta(X^{\prime})$ [22] where each observation is independent, or (cluster 2) not observed. Note that virtual evidence provides greater flexibility in modelling uncertainty compared to hard evidence where a state is observed or not. Dependent cluster mapping is necessary to optimise over all possible combinations of observed nodes and their interdependencies, which arise from propagation of evidence for inferring marginal probabilities for each scenario $p_{ijk}$ (20). DCMAP can approximate this marginal probability $p_{ijk}$ for every node $X_{i}$ via backwards propagation with evidence to layer $l$ and no evidence from forward propagation:
$p_{ijk}=\prod_{\forall X^{\prime}:l(X^{\prime})\leq l}\sum_{X^{\prime}\notin X_{i}}P\bigl{(}X^{\prime}|Par(X^{\prime})\bigr{)}\delta(X^{\prime})P^{f}_{l+1}$ (21)
@@
Clusters have been used to improve the computational efficiency of inference, and learn conditional probabilities and BN structures [5, 18]; potentially, optimal clusters can further enhance these tasks. In addition, using cost functions such as entropy, clusters could also have a practical interpretation. In non-homogeneous Hidden Markov Models or DBNs for instance, clusters could be interpreted as homogeneous portions of a non-homogeneous system, akin to changepoint detection or model switching [24, Boys2000]. Potentially, optimisation of clusters with respect to entropy could contribute to new methods in BN learning and prediction, and is a promising area for future research that is enabled by DCMAP.
[…]
Supplementary Information
6 Supplementary Information: Super-additive Objective Functions

### v6

@@ 4. Case Study
Figure 3. Overall seagrass DBN model (Wu2018) (Fig. 3a). Nodes are ovals and arrows denote conditional dependence between a parent and child node in the same time-slice. Where an arrow is labelled with a 11, the child node is in the next time-slice. Rounded rectangles denote subnetworks. Yellow nodes relate to loss and recovery in shoot density, purple nodes to recovery, green nodes to resistance, blue nodes to environmental factors and pink for all other nodes. The environment subnetwork is not shown as it is a disjoint subnetwork of the blue nodes, which are effectively inputs to the model.
@@
Table 4. Distribution of in- and out-degree (degree and number of nodes with that degree) two time-slice DBN.
@@
4.1. Computation Cost as an Objective Function
@@
Without loss of generality, consider a discrete BN. We formulate a backwards probability measure $P^{b}_{kl}$ for computing inference akin to the rearrangement of Equation (3), which forms the basis of inference techniques such as variable elimination (Daly2011).
(18) $P^{b}_{kl}=\prod_{\forall X_{i}\in X_{kl}}\sum_{X_{i}}P\bigl(X_{i}|Par(X_{i})\bigr)\prod_{X_{kl^{\prime}}\in Child(X_{kl})}\sum_{\forall X\notin X_{kl^{\prime}}}P^{b}_{kl^{\prime}}$
$P^{b}_{kl}$ comprises propagation from child cluster-layer $P^{b}_{kl^{\prime}}$, marginalising (i.e. summing) out any nodes that are not in $\boldsymbol{X}_{kl}$, multiplication by the conditional probabilities of the nodes in $\boldsymbol{X}_{kl}$, and finally marginalisation in preparation for propagation to the next antecedent cluster-layer. It is a recursive relation and for $l^{\prime}=0$, $P^{b}_{kl^{\prime}}=1$ since a cluster-layer comprising leaf nodes does not have child nodes, by Definition 1. In the context of $c\Bigl(\boldsymbol{X}_{kl^{\prime}},u(\boldsymbol{X}_{kl^{\prime}}),Child(\boldsymbol{X}_{kl^{\prime}}),u\bigl(Child(\boldsymbol{X}_{kl^{\prime}})\bigr)\Bigr)$ (10), the limits of the products and sums incorporates the effect of cluster mappings $u$, and $P(X_{i}|Par(X_{i}))$ and $P^{b}_{kl^{\prime}}$ are defined with respect to cluster-layer and child cluster-layer nodes.
Define $\mathcal{F}()$ as a normalised count of the number of clock cycles associated with performing all the multiplications and additions, where each multiplication is assumed to have a normalised count of $1$ cycle, each addition $0.6$ and each division (i.e. ratio) $3$, based on Intel Skylake CPU latencies (Fog2022). Hence,
(19) $c\Bigl(\boldsymbol{X}_{kl^{\prime}},u(\boldsymbol{X}_{kl^{\prime}}),Child(\boldsymbol{X}_{kl^{\prime}}),u\bigl(Child(\boldsymbol{X}_{kl^{\prime}})\bigr)\Bigr)=\mathcal{F}\bigl(P^{b}_{kl}\bigr)$
4.1.1. Results and Discussion
@@
Figure 4. Cost $G$ of solutions found and the iteration of their discovery for both the one and two time-slice DBNs.
@@
Figure 5. Similarity of solutions to the set of optimal solutions and their iteration of discovery for both the one and two time-slice DBNs.
@@
5. Discussion
Clustering has been applied to improve the computational feasibility of BN learning via constraints, scoring functions (i.e. using a cost function) or both (Huang2022), and the computational feasibilitiy of inference through heuristics and local optimisation or both (Forouzan2015). However, algorithms that explicitly focus on forming optimal clusters in the context of BN inference are missing. This incurs additional complexity in the form of cost dependency, where the local cost of a cluster is dependent on the mapping of connected nodes and the global cost is typically the sum of local costs. The proposed, general optimisation of BN clusters, i.e. with arbitrary criteria, advances the field by: (i) providing a benchmark for validating approximate methods against the globally optimal solution, (ii) enabling future research and development of cluster-based methods for inference including anytime optimisation (Mateescu2010) and criteria such as entropy (Lin2020) (Section 5.2), and (iii) finding clusters of homogeneous regions in a system with changing dynamics (Wu2018).
@@ 5. Discussion
Previous BN literature has used decomposability of scoring functions, such as through additive costs, to allow for local search procedures and/or optimisation of the score function; some even use dynamic programming and graph layers (Friedman1999; Kojima2010; Lu2021). Our work builds upon the application of dynamic programming and layers in BNs to facilitate global optimisation with arbitrary non-negative cost functions and cost dependency. This latter property has not been explicitly studied in the existing BN literature.
@@
The proposed algorithm, DCMAP, is proven to find all of the least cost cluster mapping solutions through conditional optimality of layers and branching (Section 3.2.3). In addition, it is theoretically shown that an equally informed algorithm cannot find these solutions in fewer iterations. This is important as optimal clustering is NP-hard in general (Buluc2016) and cost dependency would exacerbate this complexity. For instance, a breadth first search on the same one time-slice case study is unable to find any optimal solution even with ten times more iterations than DCMAP. This arises even though the number of iterations needed is similar to DCMAP for the simple illustrative BN (Fig. 1).
5.1. Case Study Discussion
@@
We developed a case study of BN clustering where clusters were used as a means to compute (i.e. infer) the marginal probabilities of a complex systems seagrass DBN. By explicitly defining the cost function as the computational cost of inference given a cluster mapping, a speed-up of 22.8 and 426 times was achieved for a one and two time-slice version of the DBN, respectively. In itself, a substantial reduction of computational complexity through clustering is useful as BN inference plays a key role in prediction and model learning, and methods for that such as Markov Chain Monte Carlo (MCMC) (Gelman2013) can require thousands of iterations. MCMC can also help better quantify uncertainty for parameter and structure learning of BNs (Daly2011) and DBNs (Shafiee2020) through the Bayesian framework. Explicit optimisation of BN clusters using DCMAP could be a key enabler for future research in BN learning and updating, especially for complex systems DBNs like our seagrass case study.
@@
Although our case study was for a discrete BN with between two and fourteen states per node with corresponding CPTs, the proposed DCMAP algorithm equally applies to discrete or continuous or hybrid BNs, DBNs, Object Oriented Bayesian Networks (OOBNs) (Koller2013) and Markov models in general, all of which can be represented as a DAG; the key adaptation required is the development of an application-specific cost function. Note that dynamic and templated models alike can be unrolled (Wu2018) into a BN and cluster mapping performed on the result. In practice, for a $k^{th}$ order dynamic model, running cluster mapping on $k$ slices is sufficient to find optimal clusters that can then be replicated over all slices.
@@
We have theoretically shown that an equally informed algorithm cannot find all optimal solutions in fewer iterations, further study is needed to better understand the computational complexity of DCMAP to find the first optimal solution and how it might vary with different BN graph structures and cost functions. Early results from the case study suggest a complexity of $O(n^{2})$ iterations to find the first least cost cluster mapping and $O(an^{2})$ iterations for all optimal solutions, where $n$ is the number of nodes and $a$ some positive integer. This aligns with previous findings of $O(n^{2})$ complexity for clustering in the absence of cost dependency (Zheng2020). However, the computational complexity could potentially vary significantly depending on the cost function employed and also the structure of the network (e.g. dense versus sparse). Additionally, it could also vary with the $\alpha$ probability for switching between best first and breadth first search (Section 3.1.2), especially for near-optimal solutions. These are all avenues for future research.
@@
5.2. Future Work: Entropy
In light of the generalisability of DCMAP given its ability to optimise for arbitrary, non-negative and dependent cost functions, one key area of future work is in the development of objective functions for specific usage scenarios. One such possibility is entropy, which for BNs is defined generally as the sum of the entropy associated with each scenario over all possible scenarios, where a scenario might be defined in terms of one or more nodes (Scutari2024; Parhizkar2018). Consider a discrete BN. The entropy associated with evidence for a given scenario indexed by $k$ is:
(20) $H_{k}=-\sum_{k=1}^{n^{k}}p_{k}\sum_{i=1}^{n}\sum_{j=1}^{n^{s}_{i}}p_{ijk}log(p_{ijk})$
@@ 5.2. Future Work: Entropy
Consider a special case of clustering where specific evidence scenarios are to be applied at one or more nodes and at one or more points in time in a DBN, with the goal of minimising global entropy. Thus, each node can be assigned to one of two clusters: (cluster 1) observed using virtual evidence $\delta(X^{\prime})$ (Pearl1988) where each observation is independent, or (cluster 2) not observed. Note that virtual evidence provides greater flexibility in modelling uncertainty compared to hard evidence where a state is observed or not. Dependent cluster mapping is necessary to optimise over all possible combinations of observed nodes and their interdependencies, which arise from propagation of evidence for inferring marginal probabilities for each scenario $p_{ijk}$ (20). DCMAP can approximate this marginal probability $p_{ijk}$ for every node $X_{i}$ via backwards propagation with evidence to layer $l$ and no evidence from forward propagation:
(21) $p_{ijk}=\prod_{\forall X^{\prime}:l(X^{\prime})\leq l}\sum_{X^{\prime}\notin X_{i}}P\bigl(X^{\prime}|Par(X^{\prime})\bigr)\delta(X^{\prime})P^{f}_{l+1}$
@@
Clusters have been used to improve the computational efficiency of inference, and learn conditional probabilities and BN structures (Daly2011; Lu2021); potentially, optimal clusters can further enhance these tasks. In addition, using cost functions such as entropy, clusters could also have a practical interpretation. In non-homogeneous Hidden Markov Models or DBNs for instance, clusters could be interpreted as homogeneous portions of a non-homogeneous system, akin to changepoint detection or model switching (Shafiee2020; Boys2000). Potentially, optimisation of clusters with respect to entropy could contribute to new methods in BN learning and prediction, and is a promising area for future research that is enabled by DCMAP.
Appendix A Appendix: Simple Example Comparison of Clustering and Computation Cost
To illustrate how clustering can support DAG inference, consider an example BN (Fig. 6a) and the fundamental task of finding posterior distributions, specifically $P(X,\boldsymbol{E}),\forall X\in\boldsymbol{X}$, Equation (3).
FGDEBAC Layer 0 Layer 1 Layer 2
(a) DCMAP Clusters
AFABDDEGCEADE
(b) Jointree Clusters
GEDCBAF
(c) Buckets
Figure 6. Example network showing nodes, directed edges, and layers (Fig. 6a). Shading denotes clusters (Fig. 6a) obtained using DCMAP, whereas the jointree (Lauritzen1988) comprises clusters (circles) and sepsets (squares) (Fig. 6b), and Bucket Elimination (Dechter1999) uses buckets (Fig. 6c). Assume all nodes have binary states.
Assuming binary states for all nodes, the Conditional Probability Table (CPT) for each node with no parents (nodes $A$, $B$, $C$) is a $2\times 1$ matrix, e.g. $P(A)=[p_{a},p_{\bar{a}}]^{T}$ for states $A=a$ and $A=\bar{a}$. For $G$ and $F$ the CPT is a $2\times 2$ matrix, e.g. $P(F|A)=\begin{bmatrix}p_{af}&p_{\bar{a}f}\\ p_{a\bar{f}}&p_{\bar{a}\bar{f}}\end{bmatrix}$ where $p_{af}$ is the probability of state $F=f$ given state $A=a$. $P(D|A,B)$ is a $2\times 4$ matrix. Multiplication of CPTs (1) containing different nodes, i.e. variables, grows the dimensionality of the resultant matrix with associated exponential growth in computation; in the discrete case, the resultant matrix contains all possible combinations of node states. We ignore evidence terms for notational simplicity, noting that evidence can be easily incorporated using $\delta(X_{j})$ with node $X_{j}$ (1). Using (3), the posterior probability $P(F)$ is:
(22) $\footnotesize P(F)$ $=\sum\limits_{A,B,C,D,E,G}{P(A)P(B)P(C)P(D|A,B)P(E|C)P(F|A)P(G|D,E)}$
(23) $P(F)$ $=\sum\limits_{A}{P(F|A)P(A)}\sum\limits_{B}{P(B)}\sum\limits_{C}{P(C)}\sum\limits_{D}{P(D|A,B)}\sum\limits_{E}{P(E|C)}\sum\limits_{G}{P(G|D,E)}$
The computation of (22) produces a matrix of $7$ dimensions, one for each node $A$ through $F$, with corresponding exponential growth in memory ($2^{7}$ joint probability values) and computation associated with multiplication and marginalisation, aka product and sum. In contrast, the maximum dimensionality of (23) obtained through Bucket Elimination (BE) (Dechter1999) is $4$, occurring at bucket $D$ (Fig. 6c). Intuitively, (22) can be rearranged, because of commutativity, to minimise computation by marginalising (summing) out nodes as early as possible (23). BE in effect employs dynamic programming to optimally compute the posterior marginal probability of a given target node, specified as part of a node ordering. In Fig. 6c, the specified order is (top to bottom) $G$, $E$, $D$, $C$, $B$, AA, $F$.
Assume that the addition of two numbers has a computational cost of $0.6$ compared to $1$ for multiplication and $3$ for divison (i.e. ratio), based on floating point operation latencies for the Intel Skylake CPU as an example (Fog2022)). Consider the computation of $\lambda_{G}(D,E)=\sum_{G}{P(G|D,E)}$ in bucket $G$ of Fig. 6c, which is the last term in (23). Nominally, this comprises multiplication of $1$ by $P(G|D,E)$ which, given every possible combination of binary states, requires $2^{3}=8$ multiplication operations and $4$ addition operations (one per combination of $D$ and $E$ states). Thus, the output $\lambda_{G}(D,E)$ is computed at a cost of $8+6\times 0.4=10.4$, and only has two dimensions (i.e. $D$ and $E$) and a joint probability distribution size of $2^{2}$. Bucket $E$ computes $\lambda_{E}(C,D)=\sum_{E}{P(E|C)}\lambda_{G}(D,E)$, requiring $2^{2}\times 2$ multiplication operations since there is an overlap in the $E$ dimension between $P(E|C)$ and $\lambda_{G}(D,E)$. Marginalisation again needs $4$ addition operations. The total computational cost is 64.4 to find $P(F)$ (23); finding the posterior for all seven nodes would incur a cost around 448, depending on input node ordering for BE.
As might be expected, other exact inference methods including Pearl’s belief propagation and polytrees, and (Lauritzen1988)’s widely-used clustering algorithm share similarities with BE as all evaluate (3) exactly (Dechter1999). However, clustering (Lauritzen1988) has a distinct advantage as it does not require a pre-specified node ordering and has less computational cost compared to BE when finding the posterior $P(X,\boldsymbol{E})$ for every node. This is achieved by performing inference in two steps: (i) forward-backward propagation along the jointree (Fig. 6b), and (ii) marginalisation of clusters to get $P(X,\boldsymbol{E})$; note that the posterior distribution given evidence $P(X|\boldsymbol{E})$ is obtained by normalising $P(X,\boldsymbol{E})$. The jointree, comprising clusters connected by sepsets, is derived from the DAG (Fig. 6a) using a heuristic process known as triangulation. forward-backward propagation involves: (a) projection from one cluster to a connected cluster via the sepset, and (b) absorption of the ratio of the new to old sepset into the receiving cluster. For example, projection of cluster $AF$ to $ABD$ (Fig. 6b) involves computing $\phi_{A}=\sum_{A}{\phi_{AF}}$, $\phi_{AF}=P(F|A)P(A)$ comprising one nominal multiplication by $P(A)$ then another by $P(F|A)$, then marginalisation for a cost of $2+4+2\times 0.6=7.2$. Absorption into $ABD$ updates $\phi_{ABD}=\phi_{ABD}\frac{\phi_{A}}{\phi^{\textrm{old}}_{A}}$, $\phi_{ABD}=P(D|A,B)$ incurring a ratio and multiplication cost of $2\times 3+8=14$. After forward-backward propagation, the posterior for $A$ is obtained from cluster $AF$ by $\sum_{F}{\phi_{AF}}$ for a cost of $2\times 0.6=1.2$; overall the cost to infer all nodes is just 132.8, which is substantially less than BE.
The above examples of BE and clustering (Lauritzen1988) show how cluster-like structures (e.g. buckets, clusters, sepsets) can facilitate more computationally efficient inference in a DAG; note that these structures differ from our definition of a cluster (Section 1). Computational efficiency is critical for MCMC-based inference (Gelman2013), as there is a need to run potentially thousands or hundreds of thousands of iterations of the models. Importantly, MCMC can help better quantify uncertainty for parameter and structure learning of BNs (Daly2011) and DBNs (Shafiee2020). Potentially, explicit optimisation of clusters using criteria associated with inference, represented as a cost, could further improve computational efficiency, make tractable MCMC for complex models, or facilitate new methods for inference such as finding homogeneous portions of a non-homogeneous system. However, this is a non-trivial task as the local cost of a cluster is dependent on the cluster mapping of connected nodes.
Appendix B Appendix: Super-additive Objective Functions